# X-ray2CTPA: Generating 3D CTPA scans from 2D X-ray conditioning

### v3

@@ Abstract
Chest X-rays or chest radiography (CXR), commonly used for medical diagnostics, typically enables limited imaging compared to computed tomography (CT) scans, which offer more detailed and accurate three-dimensional data, particularly contrast-enhanced scans like CT Pulmonary Angiography (CTPA). However, CT scans entail higher costs, greater radiation exposure, and are less accessible than CXRs. In this work we explore cross-modal translation from a 2D low contrast-resolution X-ray input to a 3D high contrast and spatial-resolution CTPA scan. Driven by recent advances in generative AI, we introduce a novel diffusion-based approach to this task. We evaluate the model’s performance using both quantitative metrics and qualitative feedback from radiologists, ensuring diagnostic relevance of the generated images. Furthermore, we employ the synthesized 3D images in a classification framework and show improved AUC in a PE categorization task, using the initial CXR input. The proposed method is generalizable and capable of performing additional cross-modality translations in medical imaging. It may pave the way for more accessible and cost-effective advanced diagnostic tools. The code for this project is available: https://github.com/NoaCahan/X-ray2CTPA.
@@ Introduction
Chest X-rays or chest radiography (CXR) are pivotal in emergency diagnostics due to their speed and accessibility, offering a cost-effective, non-invasive solution [1]. However, their 2D nature and lower soft tissue contrast limit their diagnostic ability compared to CT scans, which provide comprehensive 3D views of the various tissues including the lungs, bone and soft tissue, albeit at higher costs and radiation exposure [2, 3]. CTPA scans, a specialized form of CT with fast contrast material injection, further enhance blood vessels but may be less suitable for certain groups, including those with severe contrast allergies, or renal failure. In this work, we wish to demonstrate the use of advanced generative AI schemes for cross-model translation. In our main focus scenario, we build a system to translate 2D X-ray images into 3D CTPA scans. We explore the quality of the generated scans and their promise for earlier detection of findings in the initial CXR data, specifically focusing on Pulmonary Embolism (PE) classification. Finally, We show the generalization of the developed cross-model translation scheme using the LIDC dataset, demonstrating possible transition from CXR to synthetic thoracic CT images.
We address the classification of pulmonary embolism (PE) as present or absent, from CXR imaging as a case study in order to demonstrate the potential of transferring CXR to CT. PE is a life-threatening condition where a blood clot that develops in a limb vein travels to the pulmonary arteries where it suddenly blocks the blood flow. The gold standard for PE diagnosis is CTPA – a 3D lung CT with contrast injection. However, the use of CT is associated with exposure to ionizing radiation, and iodinated contrast, requires expensive scanners and qualified staff to operate them, and radiologists to interpret. The number of CTPAs performed continues to rise, with only 5-15% of them being positive for PE diagnosis [4]. Therefore, there is a need in the modern health care system to develop a simple, easy, fast and objective tool for diagnosing this common but life-threatening disease of PE, based on low-radiation-low-cost data, which is routinely collected for the investigation of patients with cardio-respiratory complaints utilizing advanced AI-based tools. In the emergency department, almost every patient with cardiac or respiratory issues undergoes a CXR, despite its lower tissue resolution and limited information for PE diagnosis. This underscores the opportunity to enhance the utility of CXR imaging in PE detection through innovative approaches.
@@
    Innovative Cross-Modal Translation. We introduce a novel method, leveraging the power of diffusion models, CLIP Vision encoder and adversarial guidance to translate 2D CXR images into 3D CTPA scans. This method maintains high fidelity and diagnostic relevance, as confirmed by both quantitative metrics and qualitative feedback from radiologists. To the best of our knowledge, this is the first research that uses real CXR and CTPA paired data for the application of CXR to CTPA translation.
@@
Related work
3D Image generation diffusion models in medical imaging. Image generation is one of the primary objectives of diffusion models, which has been widely applied in a variety of styles, including generating synthetic 3D medical images [10]. Medical Diffusion [11] excelled in creating high-quality medical 3D data in an unconditional setting. There are also many 3D conditional models. We list here a few but there are many more: DISPR [12] involves reconstructing 3D cells from 2D cell images and their corresponding segmentation mask, Med-DDPM model [13] is a conditional diffusion model for semantic 3D medical image synthesis. Make-A-Volume [14] leverages latent diffusion models for cross-modality 3D brain MRI synthesis and GenerateCT [15] introduced a novel approach that combines transformers and diffusion models for generating CT volumes conditioned on free-form medical text prompts. Our work represents a departure from existing works, where conditioning often relies on text prompts trained on extensive datasets or involves adding directly correlated information, such as segmentation masks or structural guides, to the generated image. In contrast, in our research the conditioning is on a different modality altogether.
CXR to CT generation. CXR to CT conversion is an under-explored area of research and there are not many works that attempt to solve this problem. X2CT-GAN [8] and it’s extension - XprospeCT[16], explores the use of Generative Adversarial Networks (GANs) for reconstructing 3D CT volumes from biplanar 2D CXRs. MedNeRF [9] introduces a deep learning model for reconstructing CT projections from as few as a single-view CXR. This approach utilizes a neural radiance field (NeRF) architecture to learn a continuous representation of CT scans. The model effectively disentangles the shape and volumetric depth of surface and internal anatomical structures from 2D images. For both papers, the CXRs used are synthetic constructions from CTs called DRR, which are similar but not identical to real CXRs. This is due to the lack of a publicly available dataset of paired CXR and CT. We argue that the task of converting DRRs to CT scans is less complex than our task because DRRs inherently possess more information than CXRs. Also, DRRs are naturally aligned and registered with CTPAs due to the way they are constructed.
@@ Results
The diffusion models were pretrained on the RSPECT dataset [18]. Then, the diffusion model was fine-tuned on our dataset in a two-stage setting. First, the model was fine-tuned in an unconditional manner, without any CXR conditioning. In the second stage, the CXR conditioning was added, and the model was further trained. Even though the datasets were relatively small, we found that each model converged and generated realistic synthetic images without a lot of hyperparameter finetuning. We present below various quantitative and qualitative evaluations of our proposed approach.
@@
Quantitative Results
To assess sample quality, we conduct quantitative evaluations utilizing various metrics. While these metrics are commonly employed and generally align with human judgment, they are not flawless indicators. There is ongoing exploration of improved metrics in evaluating sample quality, which remains an open challenge. In essence, appraising the visual accuracy of synthetic images, particularly within medical imaging, continues to be a complex task where small details may hold significant importance [19]. Keeping in mind these limitations, we evaluated the quality of the generated CTPA volumes using four metrics:
@@
    LPIPS [24]: LPIPS essentially computes the similarity between the activations of two image patches for some pre-defined neural network. This measure has been shown to match human perception well. A low LPIPS score indicates that image patches are perceptually similar.
We computed these metrics by evaluating the generated samples against their corresponding ground truth samples in our test set. Table 1 presents the quantitative generation results for the CTPA dataset.
@@
Qualitative Results
We conducted a short expert evaluation study, with the generated scans evaluated by two expert Radiologists. First, experts were presented with 20 CTPA volumes, comprising an equal mix of 10 real and 10 synthetic ones. Readers were tasked with assessing the visual quality of these volumes and determining whether each volume was real or synthetic. Both readers accurately distinguished between most synthetic and real volumes, citing the inferior image quality of the synthetic scans compared to the real scans as the primary reason. As a followup, we asked the experts to rate the synthetic images in terms of anatomic correctness and slice consistency. The readers were asked to rate 10 generated images from the test set. All readers rated the scans as mostly realistic with only minor unrealistic areas. In addition, the images were considered consistent across slices with most images assessed as exhibiting only minor or no anatomic inconsistencies. We conclude from this initial study that our architecture can generate synthetic images that appear somewhat realistic to radiologists but further work needs to be done if we wish to achieve realistic-looking images. We hypothesize that better-quality output images can be generated once more data is used in the training phase. The small inconsistencies may be mitigated by adding a super-resolution stage, increasing the sample size, or using a pre-trained diffusion model and adapting it to our problem using tools such as ControlNet [25].
Ablation study on model architecture choices
@@
Table 2 evaluates several design choices and the contribution of different network components. We examined different loss functions as alternatives for the model training. In addition to $L_{1}$ loss which minimizes the pixel distance between the generated sample and the ground truth, we compared the use of adding perceptual loss [24], adding adversarial loss using a discriminator and classification loss. As can be seen from the results, adding the perceptual loss improved the results compared to using only $L_{1}$ loss. However, while the addition of adversarial and classification loss did not enhance the quantitative metrics, they both played important roles for our model. Adding the adversarial loss resulted in more visually realistic samples. The classification loss improved the classification results. This is presented in Fig. 4. Additionally, we modified the pretraining process to include the RSPECT dataset not only in as an unconditional dataset for CTPA scans but also in a conditional setting by incorporating synthetic CXR images or digitally reconstructed Radiographs (DRRs) created from the CTPA scans. Adding this pretraining step highly boosted our performance by all metrics.
@@
We also present the results of experiments that failed to achieve optimal outcomes. We attempted to use classifier-free guidance (CFG) for both CXR conditioning and class label conditioning, but neither approach improved the results. Additionally, we experimented with variations in sample size. The spatial resolution of 512 likely yielded sub-optimal results due to our small sample size, and the resolution of 128 failed due to the limitations of the pre-trained VAE we used, which was not trained on smaller image sizes. Lastly, we also attempted to use the pre-trained 3D-VQGAN from medical diffusion [11], which was pre-trained on the LIDC dataset, but it failed to improve results on the CTPA dataset.
@@
In addition to the quantitative results, we also compare the visual differences of using different loss functions (Fig. 4). All loss functions were trained as a weighted combination with the $l1$ loss. In general, it is evident that the generated samples are like the ground truth, although not identical. Supporting the quantitative results, the LPIPS loss improves both the visual appearances of the generated samples and seem to guide the scans to be closer in structure and color contrast to the ground truth. The pretraining step using the DRRs as conditioning further enhances the LPIPS loss improvements and provides the most realistic images which are also the most close to the ground truth. The classification loss adds distortion to the samples, which is expected as its purpose is only to aid the classification task. The adversarial loss provides very realistic images but, in some cases, the generated sample is derived further away from the ground truth. The weighted combination of all losses (Fig. 4 - "Full") provides the best tradeoff between all losses.
@@
Synthetic 3D CTPAs can be used to enhance PE classification
As part of our evaluation, we aimed to assess the generated CTPA’s suitability to be used for a classification task. Specifically, we applied the generated data to the task of PE classification. As mentioned in the Introductory section, CTPA is the gold standard for PE diagnosis, and PE is usually not noticeable on CXRs due to their non-contrast enhanced and 2D nature. We hypothesis that by training the diffusion model with matching CXR and CTPA pairs, an alignment is constructed between the two modalities that can stir the generated 3D-image so that it can improve the classification results compared to a classification using CXRs. Table 3 presents the results for this classification task.
The classification of PE using unimodal data: CTPA only and CXR only data, are used as our baselines. The results present our sixfold cross-validation analysis, by presenting the average AUC, with a 95% confidence interval, calculated across all folds using the Delong method [26]. We set our operating point based on the Youden’s J-Score statistic that maximizes the sum of sensitivity and specificity on the validation set. For the numerical threshold that separates the predicted classes, we utilized the standard definition of the operating point.
As expected, the real CTPA classification results outperform the CXR results achieving an AUC of 0.827 [95% CI: 0.75-0.92] compared to 0.691 [95% CI: 0.54-0.77]. The classification using the generated CTPAs significantly outperformed the CXR classification results with an AUC of 0.803 [95% CI: 0.72-0.9]. In addition, the specificity and NPV of the model improve as well compared to the CXR classifier. In a clinical context, this outcome suggests a higher number of true negative results, leading to a reduction in unnecessary CTPA scans performed. It is evident that using the synthetic CTPAs enhance performance compared to using solely CXRs. For the CTPAs classification, we used the latent output representation from the 2D-VAE used in our diffusion model. The CXRs were classified using RadImageNet [27], a ResNet-50 based model pretrained on millions of radiologic images.
Model AUC Accuracy Specificity Sensitivity PPV NPV
Baselines
CTPA only classifier $0.827$ [$0.75$-$0.92$] ±$\pm 0.05$ $76.92$ $78.18$ $73.91$ $58.62$ $87.76$
Generated Data
All results are statistically significant with Kolmogorov–Smirnov test, with $p\leq 0.05$
Generalizability
We wanted to evaluate the ability of our model to generate other modalities in a similar setting. To do so, we fine-tuned our model on an additional dataset. We used chest CT studies from the publicly available Lung Image Database Consortium (LIDC) dataset [28]. This dataset contains 1010 (thoracic CT) individual studies. We noticed that in addition to the available CT scans, around 290 scans also included matching CXRs. We fine-tuned the model using low rank adaptation (LoRA) [29].
@@
Failed examples
While X-ray2CTPA model can generate high-quality 3D CTPA samples that closely resemble the ground truth in structure and appearance, there are instances where it slightly diverges or fails to accurately replicate the ground truth. Some examples from our test set as presented in Fig. 6. The main areas where the model fails are around small veins or when rare abnormalities or additional pathologies (other than PE) are presented. In some of these cases the model generates the pathology in some variant but may replicate it in a wrong location or fail to generate the pathology whatsoever. We hypothesis that this issue may be caused by the high variability nature of the CTPA samples and due to the fact, that in many cases PE is secondary to other patient co-morbidities. We believe that training the model with a larger dataset may solve this issue in future work.

### v4

# X-ray2CTPA: Leveraging Diffusion Models to Enhance Pulmonary Embolism Classification

@@ Abstract
Chest X-rays or chest radiography (CXR), commonly used for medical diagnostics, typically enables limited imaging compared to computed tomography (CT) scans, which offer more detailed and accurate three-dimensional data, particularly contrast-enhanced scans like CT Pulmonary Angiography (CTPA). However, CT scans entail higher costs, greater radiation exposure, and are less accessible than CXRs. In this work, we explore cross-modal translation from a 2D low contrast-resolution X-ray input to a 3D high contrast and spatial-resolution CTPA scan. Driven by recent advances in generative AI, we introduce a novel diffusion-based approach to this task. We employ the synthesized 3D images in a classification framework and show improved AUC in a Pulmonary Embolism (PE) categorization task, using the initial CXR input. Furthermore, we evaluate the model’s performance using quantitative metrics, ensuring diagnostic relevance of the generated images. The proposed method is generalizable and capable of performing additional cross-modality translations in medical imaging. It may pave the way for more accessible and cost-effective advanced diagnostic tools. The code for this project is available: https://github.com/NoaCahan/X-ray2CTPA.
@@ Introduction
CXRs are pivotal in emergency diagnostics due to their speed and accessibility, offering a cost-effective, non-invasive solution [1]. However, their 2D nature and lower soft tissue contrast limit their diagnostic ability compared to CT scans, which provide comprehensive 3D views of the various tissues including the lungs, bone and soft tissue, albeit at higher costs and radiation exposure [2, 3]. CTPA scans, a specialized form of CT with fast contrast material injection, further enhance blood vessels but may be less suitable for certain groups, including those with severe contrast allergies, or renal failure. In this work, we wish to demonstrate the use of advanced generative AI schemes for cross-model translation and to use the generated data to enhance a classification task. In our main focus scenario, we build a system to translate 2D chest X-ray images into 3D CTPA scans. We explore the quality of the generated scans and their promise for earlier detection of findings in the initial CXR data, specifically focusing on PE classification. Finally, We show the generalization of the developed cross-model translation scheme using the LIDC dataset, demonstrating possible transition from CXR to synthetic thoracic CT images.
We address the classification of PE as present or absent, from CXR imaging as a case study in order to demonstrate the potential of transferring CXR to CT. PE is a life-threatening condition where a blood clot that develops in a limb vein travels to the pulmonary arteries where it suddenly blocks the blood flow. The reference standard for PE diagnosis is CTPA – a 3D lung CT with contrast injection. However, the use of CT is associated with exposure to ionizing radiation, and iodinated contrast, requires expensive scanners and qualified staff to operate them, and radiologists to interpret. The number of CTPAs performed continues to rise, with only 5-15% of them being positive for PE diagnosis [4]. Therefore, there is a need in the modern health care system to develop a simple, easy, fast and objective tool for diagnosing this common but life-threatening disease of PE, based on low-radiation-low-cost data, which is routinely collected for the investigation of patients with cardio-respiratory complaints. In the emergency department, almost every patient with cardiac or respiratory issues undergoes a CXR, despite its lower tissue resolution and limited information for PE diagnosis. This underscores the opportunity to enhance the utility of CXR imaging in PE detection through innovative approaches.
Given the substantial drawbacks associated with CT - such as high costs, significant radiation exposure, and logistic bottlenecks - there is a critical need to leverage more accessible data sources like CXRs. However, CXRs alone often lack sufficient detail for definitive diagnosis, underscoring the opportunity to introduce advanced AI-based methods. By harnessing generative and discriminative deep learning schemes, we aim to supplement the limited information in CXRs with high-fidelity, synthesized 3D CTPA representations that could ultimately improve diagnostic outcomes and reduce unnecessary scans.
@@
In general, CXR to CT conversion is an under-explored area of research and there are not many works that attempt to solve this problem. X2CT-GAN [8] and it’s extension - XprospeCT[16], explores the use of GANs for reconstructing 3D CT volumes from biplanar 2D CXRs. MedNeRF [9] introduces a deep learning model for reconstructing CT projections from as few as a single-view CXR. This approach utilizes a neural radiance field (NeRF) architecture to learn a continuous representation of CT scans. The model effectively disentangles the shape and volumetric depth of surface and internal anatomical structures from 2D images. For both papers, the CXRs used are synthetic constructions from CTs called DRR, which are similar but not identical to real CXRs. This is due to the lack of a publicly available dataset of paired CXR and CT. We argue that the task of converting DRRs to CT scans is less complex than our task because DRRs inherently possess more information than CXRs. Also, DRRs are naturally aligned and registered with CTPAs due to the way they are constructed.
Image generation is one of the primary objectives of diffusion models, which has been widely applied in a variety of styles, including generating synthetic 3D medical images [10]. Medical Diffusion [11] excelled in creating high-quality medical 3D data in an unconditional setting. There are also many 3D conditional models. We list here a few but there are many more: DISPR [12] involves reconstructing 3D cells from 2D cell images and their corresponding segmentation mask, Med-DDPM model [13] is a conditional diffusion model for semantic 3D medical image synthesis. Make-A-Volume [14] leverages latent diffusion models for cross-modality 3D brain MRI synthesis and GenerateCT [15] introduced a novel approach that combines transformers and diffusion models for generating CT volumes conditioned on free-form medical text prompts.
Our work represents a departure from existing works, where conditioning often relies on text prompts trained on extensive datasets or involves adding directly correlated information, such as segmentation masks or structural guides, to the generated image. In contrast, in our research the conditioning is on a different modality altogether.
@@
    Innovative Cross-Modal Translation. We introduce a novel method, leveraging the power of diffusion models, CLIP Vision encoder, and adversarial guidance to translate 2D CXR images into 3D CTPA scans. This method maintains high fidelity and diagnostic relevance, as demonstrated by both quantitative metrics and improved PE classification performance using the generated data. To the best of our knowledge, this is the first study to utilize real paired CXR and CTPA data for CXR-to-CTPA translation.
@@ Results
We present below the results of applying our proposed generation approach for PE classification enhancement from CXR as well as various quantitative and qualitative evaluations of the generated samples. Finally, we perform an ablation study in which we explore different model architecture choices.
@@
Visualizing the Generated Data
Using Generated 3D CTPAs to Enhance PE Classification
As part of our evaluation, we aimed to assess the generated CTPA’s suitability to be used for a classification task. Specifically, we applied the generated data to the task of PE classification. As mentioned in the Introductory section, CTPA is the reference standard for PE diagnosis. PE is usually not noticeable on CXRs due to their non-contrast enhanced and 2D nature. We hypothesis that by training the diffusion model with matching CXR and CTPA pairs, an alignment is constructed between the two modalities that can stir the generated 3D-image so that it can improve the classification results compared to a classification using CXRs.
Table 1 presents the results for this classification task. We start with a baseline: The classification of PE using unimodal data, CTPA only and CXR only data. The results present our sixfold cross-validation analysis, by presenting the average AUC, with a 95% confidence interval, calculated across all folds using the Delong method [18]. We set our operating point based on the Youden’s J-Score statistic that maximizes the sum of sensitivity and specificity on the validation set. For the numerical threshold that separates the predicted classes, we utilized the standard definition of the operating point.
Table 1 presents the following findings: First, we note the baseline results, where using CTPA data achieves an AUC of 0.90 [95% CI: 0.88-0.97], whereas using CXR data yields a significantly lower AUC of 0.691 [95% CI: 0.54-0.77]. When utilizing the generated data, the model achieves an AUC of 0.803 [95% CI: 0.72-0.9]. This is a significant improvement over the original CXR data ($\alpha\leq 0.05$), approaching the performance benchmark of true CTPA data. Furthermore, the model demonstrates improvements in specificity and NPV compared to the CXR classifier. In a clinical context, this outcome suggests a higher number of true negative results, potentially reducing the number of unnecessary CTPA scans performed. These findings underscore that synthetic CTPA data enhances model performance compared to relying solely on CXRs.
Model AUC Accuracy Specificity Sensitivity PPV NPV
Baselines
CTPA only classifier $0.90$ [$0.88$-$0.97$] ±$\pm 0.04$ $84.05$ $83.86$ $82.74$ $69.78$ $92.32$
Generated Data
All AUC results are statistically significant with Delong-ROC test, with $p\leq 0.05$
Generative Model Exploration
Quantitative Results.
To assess the generated sample quality, we conduct quantitative evaluations utilizing various metrics. While these metrics are commonly employed and generally align with human judgment, they are not flawless indicators. There is ongoing exploration of improved metrics in evaluating sample quality, which remains an open challenge. In essence, appraising the visual accuracy of synthetic images, particularly within medical imaging, continues to be a complex task where small details may hold significant importance [19]. Keeping in mind these limitations, we evaluated the quality of the generated CTPA volumes using four metrics:
@@
    LPIPS [24]: perceptual similarity (LPIPS) essentially computes the similarity between the activations of two image patches for some pre-defined neural network. This measure has been shown to match human perception well. A low LPIPS score indicates that image patches are perceptually similar.
We computed these metrics by evaluating the generated samples against their corresponding ground truth samples in our test set. Table 2 presents the quantitative generation results for the CTPA dataset. We recognize that these quantitative metrics alone may not fully capture the visual quality of our generated scans. Additional comparisons and analyses are provided in the subsequent Ablation Study section.
@@
Mapping the Latent Space of Generated Samples.
In Fig. 4 we present the visualization results of the t-SNE [25] and PCA [26] of the latents generated by the conditional model compared to their corresponding ground truth samples from the test set. For comparison, we also present the t-SNE and PCA of the latents generated from an unconditional model. As can be seen from the results, the generated samples closely follow the ground-truth samples. This is especially apparent in comparison to the unconditional samples.
Figure 4: t-SNE and PCA visualizations results of generated latents compared to the ground truth samples on the test set. The ground truth samples are marked in blue, the generated samples are marked in orange and the unconditional samples in green.
Ablation study on model architecture choices.
@@
Table 3 evaluates several design choices and the contribution of different network components. We examined different loss functions as alternatives for the model training. In addition to $L_{1}$ loss which minimizes the pixel distance between the generated sample and the ground truth, we compared the use of adding perceptual loss [24], adding adversarial loss using a discriminator and classification loss. As can be seen from the results, adding the perceptual loss improved the results compared to using only $L_{1}$ loss. However, while the addition of adversarial and classification loss did not enhance the quantitative metrics, they both played important roles for our model. Adding the adversarial loss resulted in more visually realistic samples. The classification loss improved the classification results. This is presented in Fig. 5. Additionally, we modified the pretraining process to include the RSPECT dataset not only in as an unconditional dataset for CTPA scans but also in a conditional setting by incorporating synthetic CXR images or DRRs created from the CTPA scans. Adding this pretraining step highly boosted our performance by all metrics.
@@
We also present the results of experiments that failed to achieve optimal outcomes. We attempted to use classifier-free guidance (CFG) for both CXR conditioning and class label conditioning, but neither approach improved the results. Additionally, we experimented with variations in sample size. The spatial resolution of 512 likely yielded sub-optimal results due to our small sample size, and the resolution of 128 failed due to the limitations of the pretrained VAE we used, which was not trained on smaller image sizes. Lastly, we also attempted to use the pretrained 3D-VQGAN from medical diffusion [11], which was pretrained on the LIDC dataset, but it failed to improve results on the CTPA dataset.
@@
In addition to the quantitative results, we also compare the visual differences of using different loss functions (Fig. 5). All loss functions were trained as a weighted combination with the $l1$ loss. In general, it is evident that the generated samples are like the ground truth, although not identical. Supporting the quantitative results, the LPIPS loss improves both the visual appearances of the generated samples and seem to guide the scans to be closer in structure and color contrast to the ground truth. The pretraining step using the DRRs as conditioning further enhances the LPIPS loss improvements and provides the most realistic images which are also the most close to the ground truth. The classification loss adds distortion to the samples, which is expected as its purpose is only to aid the classification task. The adversarial loss provides very realistic images but, in some cases, the generated sample is derived further away from the ground truth. The weighted combination of all losses (Fig. 5 - "Full") provides the best trade-off between all losses.
@@
Exploring the use of the 2D-VAE Vs. 3D-VQGAN.
We explore two alternatives as compression components for the latent diffusion process: a 3D-VQGAN applied to the entire volume and a 2D-VAE applied to individual 2D slices of the 3D volume, which are subsequently concatenated. Both compressors were used "off-the-shelf" without additional training on our dataset. The 3D-VQGAN was sourced from the medical diffusion model [11] and trained on the LIDC dataset [27]. Table 4 presents the quantitative comparison between these two approaches. The results compare reconstructed samples from the test set, processed through each compression component, against the ground truth. The quantitative analysis shows that both methods perform well, with the 2D-VAE outperforming the 3D-VQGAN, despite not being trained on medical data. Notably, the FVD score of the 2D-VAE demonstrates that the continuity between slices within the 3D volume is preserved.
Table 4: Quantitative evaluation of the reconstructed test set samples from the 2D-VAE.
Compressing Component Metric
FVD $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
2D-VAE $0.05\pm 0.05$ $31.00\pm 2.5$ $0.84\pm 0.09$ $0.03\pm 0.02$
3D-VQGAN $0.1\pm 0.2$ $26.35\pm 2.8$ $0.74\pm 0.08$ $0.05\pm 0.03$
Fig. 6 shows consecutive 2D slices from the axial, coronal, and sagittal views of the 3D volume, comparing the ground truth to the samples reconstructed by the 2D-VAE. The results indicate that the transitions between slices remain continuous across all views, and the inter-slice changes are preserved during the compression and decompression processes of the 2D-VAE.
Figure 6: Visual results of ground truth and reconstructed samples after applying 2D-VAE compression and decompression. Presented are consecutive slices from three views of the 3D volume ((A) axial, (B) coronal and (C) sagittal)
Generalizability.
We wanted to evaluate the ability of our model to generate other modalities in a similar setting. To do so, we fine-tuned our model on an additional dataset. We used chest CT studies from the publicly available Lung Image Database Consortium (LIDC) dataset [27]. This dataset contains 1010 individual studies (thoracic CT) . We noticed that in addition to the available CT scans, around 290 scans also included matching CXRs. We fine-tuned the model using low rank adaptation (LoRA) [28].
@@
User Studies
We conducted two short expert evaluation studies, with the generated scans separately evaluated by two board certified radiologists.
Generated Image Quality.
In this study, experts were separately presented with 20 CTPA volumes, comprising an equal mix of 10 real and 10 synthetic ones. Readers were tasked with assessing the visual quality of these volumes and determining whether each volume was real or synthetic. Both readers accurately distinguished between most synthetic and real volumes, citing the inferior image quality of the synthetic scans compared to the real scans as the primary reason. As a followup, we asked the experts to rate the synthetic images in terms of anatomic correctness and slice consistency. The readers were asked to rate 10 generated images from the test set. All readers rated the scans as mostly realistic with only minor unrealistic areas. In addition, the images were considered consistent across slices with most images assessed as exhibiting only minor or no anatomic inconsistencies. We conclude from this initial study that our architecture can generate synthetic images that appear somewhat realistic to radiologists but further work needs to be done if we wish to achieve realistic-looking images. We hypothesize that better-quality output images can be generated once more data is used in the training phase. The small inconsistencies may be mitigated by adding a super-resolution stage, increasing the sample size, or using a pre-trained diffusion model and adapting it to our problem using tools such as ControlNet [29].
PE categorization from CXR data - with and without Generated CTPA data.
We compare radiologists’ performance when interpreting conventional CXRs for PE categorization versus assessing PE from the generated CTPA images. In the initial phase with CXR alone (20 cases: 10 PE, 10 non-PE), the readers struggled to identify the confirmed emboli, producing an overall diagnostic accuracy close to chance levels (approximately 50%), with very low sensitivity. This poor performance reflects the well-known limitation of chest radiography in diagnosing PE. In contrast, during the second phase, the same cases were reassessed, incorporating the corresponding generated 3D CTPA volumes. Table 6 summarizes the results of this study. Here, the radiologists correctly identified the majority of PE cases. The diagnostic accuracy improved substantially to 75% (p-value = 0.025 and 0.033 using McNemar’s test), with a clear increase in true-positive detections, demonstrating a significant boost over the CXR baseline. These results indicate that the synthesized CTPA imaging provided critical visual clues that were absent on CXRs, substantially enhancing the radiologists’ ability to detect PE.
Table 6: Expert evaluation of PE categorization task.
Experts Accuracy Sensitivity Specificity
Reader 1 $0.75$ $0.8$ $0.7$
Reader 2 $0.75$ $0.9$ $0.6$
Ablation Study on CTPA Classifier
We implemented two different classifiers for the real and synthetic CTPA scans. The first is the INSPECT classifier [30], which is implemented in the pixel space, and the second is a 3D DenseNet3D-121 classifier trained in the latent space of the diffusion model. The detailed architectural and training details for both classifiers are provided in the Methods section.
Table 7 summarizes the comparison results of the two classifiers on the original and synthesized CTPA scans. The data set in its latent form is limited by the ability of the diffusion model to generate only a maximum of 64 frames along the z-axis. Therefore, we benchmarked INSPECT with three sets of slices: 512, 128, and 64. Using all 512 slices, INSPECT achieves the strongest performance (AUC = 0.90; 95 % CI [0.88–0.97]). As the slice count decreases, its AUC drops accordingly. The results of the latent classifier and the 64-slice INSPECT model receive similar results, where INSPECT slightly outperforms the latent classifier with a gain of 2% On the synthesized scans, the latent-space classifier substantially surpasses the INSPECT classifier. We attribute this gap to the training strategy - during diffusion training, we use the relation between CXR and 3D CTPA to infuse the latent classifier with information unavailable to a model trained solely in pixel space.
Table 7: Ablation on the CTPA classifier. We compare the average AUC scores (with a 95% confidence interval) of the PE classification from the original CTPAs and generated CTPAs of our test set.
Model AUC Accuracy Specificity Sensitivity
Original CTPA data
CTPA INSPECT 512 slices $0.90$ [$0.88$-$0.97$] ±$\pm 0.08$ $84.05$ $83.86$ $82.74$
CTPA INSPECT 128 slices $0.88$ [$0.8$-$0.95$] ±$\pm 0.1$ $79.5$ $89.6$ $73.91$
CTPA INSPECT 64 slices $0.85$ [$0.77$-$0.92$] ±$\pm 0.08$ $82.3$ $84.7$ $74.9$
CTPA latent $0.827$ [$0.75$-$0.92$] ±$\pm 0.05$ $76.92$ $78.18$ $73.91$
Generated Data
Generated CTPA latent classifier $0.80$ [$0.72$-$0.9$] ±$\pm 0.07$ $74.0$ $74.1$ $73.91$
Generated CTPA INSPECT classifier $0.74$ [$0.68$-$0.82$] ±$\pm 0.06$ $72.1$ $71.88$ $71.9$
PE Classifier Grad-CAM Visualization Results
We employed gradient back-propagation techniques for model interpretability. We highlighted the features selected by the model during prediction and identified locations and slices in the 3D-CTPA imaging scan that contributed the most to the classification using 3D gradient-weighted class activation mapping (Grad-CAMs) [31]. For our implementation we used Captum open source package, an extensible library for model interpretability built on PyTorch [32]. Fig. 8 illustrates the heatmap visualizations generated by back-propagating through the 3D PE classification model - highlighting areas that most influenced the model’s decision. We anticipated that the model would focus on regions associated with PE as well as descriptors correlated with high-risk PE, such as the heart region and hepatic veins. As expected, the network concentrated on areas containing PE clots and the heart (particularly the right and left ventricular (RV, LV) chambers). We note that the Grad-CAM predictions cover a broader region around the PE clots, resulting in less precise localization. We hypothesis this is due to performing the classification in the latent space, where spatial resolution is naturally lower.
[…]